Roger has $2.60 in quarters and dimes. He has 14 coins altogether. How many coins of each kind does he have?
Answer by ikleyn(52817)

.
Roger has $2.60 in quarters and dimes. He has 14 coins altogether.
How many coins of each kind does he have?
~~~~~~~~~~~~~~~

Q quarters;  (14-Q) dimes.


Write the total money equation

    25Q + 10*(14-Q) = 260   cents.


Simplify and find Q

    25Q + 140 - 10Q = 260

    25Q - 10Q = 260 - 140

        15Q   =    120

          Q   =    120/15 = 8.


ANSWER.  8 quarters and (14-8) = 6 dimes.

Solved.

------------------

For introductory coin problems,  look into the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Solving coin problems mentally by grouping without using equations
in this site.

Read them attentively and become an expert in this field.